OpenStudy (anonymous):

can someone plz check my answer What is the denominator of the simplified form of x^3/6/3x i got 6x

OpenStudy (anonymous):

i think it is wrong

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can you explain to me how to do it

OpenStudy (anonymous):

(x^3/6)/3x=x^3/6*3*x =x^2/18

OpenStudy (anonymous):

oh ok i see and i just simplify that

OpenStudy (anonymous):

good

OpenStudy (anonymous):

ok so it would be x=9 but that isnt one of my choices

OpenStudy (anonymous):

my choices are a.1 b.x^2/6 c.6x d.6/x^2

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

do u mean this:(x^3/6)/3x ?

OpenStudy (anonymous):



OpenStudy (anonymous):

that is the problem sorry for the bad drawing

OpenStudy (anonymous):

x*3*x/3*6=x^2/6

OpenStudy (anonymous):

@daddyssweetbaby ,ok?

OpenStudy (anonymous):

ok yes i see that i understand

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

if i post another one can you check and see if my answer is right